A farmer buys 100 animals for $100 total. The animals include at least 1 cow, 1 pig and 1 chicken. A cow costs $10, a pig costs $3 and a chicken costs $0.50. If the farmer buys 5 times as many cows as pigs, how many of each animal does he buy?


Answer by ikleyn(52778)

.
A farmer buys 100 animals for $100 total. The animals include at least 1 cow, 1 pig and 1 chicken.
A cow costs $10, a pig costs $3 and a chicken costs $0.50. If the farmer buys 5 times as many cows as pigs, how many of each animal does he buy?
~~~~~~~~~~~~~~~~


Let x be the number of pigs.

Then the number of cows is 5x  and the number of chicken is 100-x-5x = 100-6x.


Then the total money equation is


    10*(5x) + 3*x + 0.50*(100-6x) = 100  dollars.


Simplify and solve


    50x + 3x + 50 - 3x = 100

    50x = 50

    x = 50/50 = 1 is the number of pigs.


ANSWER.  1 pig, 5 cows and 100-1-5 = 94 chicken.

Solved.

A miracle is that this problem can be solved using one single equation in one single unknown.